# Tree balance in phylogenetic models

### v3

@@ 4 Results
In this section, we summarize the most important results obtained in our simulation study. Figures 37 show the power (with 95%-confidence bands) of the various TSS to correctly identify trees generated under models distinct from the Yule model as not having been generated under the Yule model. The power is plotted either against the leaf number $n$ (Figure 3) or against a tree model parameter (Figures 47). Note that all figures follow the legend of Figure 3. Further result figures can be found in Section 5 of the supplementary material [40]. As evident from all our figures, the curves look quite different for different alternative models. In the following, we study some aspects of these figures and conclusions to be drawn more in-depth.
“Good” and “bad” TSS
@@
A central goal of our study was to identify TSS that exhibit a high power for discriminating between different (macro)evolutionary models. Our results show that that the performance of different TSS clearly depends on the alternative models and also on the null model used (recall that we use the Yule model as the null model throughout the present manuscript, but in Section 6 of the supplementary material [40] we also briefly discuss other null models). Even for a fixed alternative model, there are sometimes striking differences for different parameter choices. For example, considering the log-Brownian models ins Figure 6, we can observe that the $B_{2}$ index and the equal weights Colless index are among the worst TSS for small values of $\sigma_{\lambda}$, whereas they are among the most powerful TSS for larger values of $\sigma_{\lambda}$. A related observation is that in some cases, there are few to no crossings between the power curves for different TSS (e.g., Figures 4 and 5), implying that the ranking of the TSS by their power does not change across the different settings tested, whereas for others, there are significant changes (e.g., Figures 5 and 6).
@@
Nevertheless, we can observe that there are some TSS that (almost) consistently outperform the others, such as the $\widehat{s}$-shape statistic, whereas others, such as the modified maximum difference in width or the rooted quartet index, tend to perform rather poorly. In all cases, there are notable exceptions, though. In the case of the $\widehat{s}$-shape statistic it is interesting to note that it is outperformed by other TSS (such as the $B_{2}$ index or the Sackin index) in the symmetric simple Brownian model (Figure 5) as well as for parameter-model combinations close to the Yule model in the asymmetric log-Brownian model (Figure 6)). In case of the modified maximum difference in widths, exceptions to its overall relatively poor performance are for instance Aldous’ $\beta$ splitting model (Figure 4) and the biased speciation model (Figure 5). Interestingly, the biased speciation model is also an exception to its overall relatively poor performance. Another exception for the rooted quartet index is the simple Brownian model (Figure 5). It is also noticeable that while the power of TSS should increase with increasing $n$, and this is indeed the case for almost all TSS (see Figure  3 and the differences between $n=30$ and $n=100$ in Figure 6), the power of the modified maximum difference in widths does not seem to follow this pattern for the PDA model (Figure 3).
Additionally, it is interesting to note that Fischer et al. [78, Chapter 26] recently observed that the $\widehat{s}$-shape statistic had a high resolution for two empirical data sets (meaning that it showed few ties), whereas the modified maximum difference in widths showed a low resolution (meaning that it showed many ties). This could indicate a connection between resolution and power for some but not all TSS. For the rooted quartet index, for instance, Fischer et al. [78, Chapter 26] observed a high resolution while we observe a relatively low power of the rooted quartet index in the present analysis. Nevertheless, exploring why certain TSS behave in distinct ways and whether there is a connection between power and resolution for some TSS is an interesting direction for future research.
@@
Finally, we remark that it is often the case that several TSS perform almost equally well or equally poorly, indicated by their curves being clumped together, as is for instance the case for the top curves in the simple Brownian model (Figure 5). In other cases, the curves are clearly separated, as for instance in the biased speciation model (also Figure 5). Interestingly, the total cophenetic index and the quadratic Colless index show extremely similar behavior (their lines are nearly always right on top of each other), and it would be interesting to investigate this further. In summary, we thus reiterate that what is the best and worst TSS depends on the model setting, highlighting the importance of performing a power analysis such as the one presented here prior to downstream analyses.
@@
For parameter settings that correspond to the Yule model, the power of all TSS should be close to 0.05 and this is clearly visible in the corresponding figures (e.g., for $\beta=0$ or $\alpha=0$ in Figure 4, or $\mu=0$ in Figure 7). Interestingly, we also observe a power close to 0.05 for almost all TSS for the biased speciation model for $\zeta$ between 0.20 and 0.25 (Figure 5). This could indicate that the biased speciation model is very close to or coincides with the Yule model for $\zeta$ in this range. However, it could also mean that none of the TSS employed in this study can detect deviation from the Yule model in this setting, even though the models might theoretically not be identical. This leads us to the following observation for fitting tree models to empirical data: When a collection of empirical trees is tested against a specific evolutionary model and a low power is observed for all TSS, this can be an indication of a good fit of the model to the data. However, low power is only a sufficient and not a necessary condition for a good model fit, as it could be the case that other TSS are needed to detect differences between the data and the model. Thus, caution is warranted. A good example illustrating this can be found in Figure 20 in the supplementary material [40], where we use Aldous’ $\beta$ splitting model with $\beta=-1$ as the null model and the alternative birth-death model as the alternative model. In this case, almost none of the TSS can tell the alternative birth-death trees with parameter $\mu=0.4$ apart from the Aldous’ $\beta=-1$ trees, except for the $B_{2}$ index and the Furnas rank (where the latter is even more surprising as it is primarily used as a tree enumeration method and not often employed in applications). Without including these two indices, one might have concluded that both models are equivalent.
@@
Figure 3: The power of all TSS to correctly identify trees generated under the PDA model as not having been generated under the Yule model as a function of $n$. Note that for $n=10$ and $n=15$ the power values are based on the exact TSS distribution under Yule.
Figure 4: The power of all TSS to correctly identify trees generated under the Aldous’ $\beta$ splitting model (left) and Ford’s $\alpha$ model (right) as not having been generated under the Yule model as a function of $\beta$ (resp. $\alpha$). Since for $\beta=0$, Aldous’ splitting model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case (as specified with the level of significance). Similarly, since for $\alpha=0$, Ford’s $\alpha$ model corresponds to the Yule model, all TSS rejected $\approx 0.05$ of the trees in that case.
Figure 5: The power of all TSS to correctly identify trees generated under the biased speciation model (left) and the asymmetric simple Brownian model (right) as not having been generated under the Yule model as a function of $\zeta$ (resp. $\sigma$).
Figure 6: The power of all TSS to correctly identify trees generated under the symmetric (left) and asymmetric (right) punctuated(-intermittent) log-Brownian model as not having been generated under the Yule model as a function of $\sigma_{\lambda}$. The starting trait value is $x_{0}=10$ with constant $\sigma_{x}=1$.
Figure 7: The power of all TSS to correctly identify trees generated under the alternative birth-death model as not having been generated under the Yule model as a function of $\mu$ for two different different tree sizes ($n=30$ and $n=100$).
@@ 5 Conclusion and outlook
In most of our analyses, the $\widehat{s}$-shape statistic was the most powerful statistic discriminating alternative models from the Yule model. While this nicely complements a study by Blum at al. [27], who found that the $\widehat{s}$-shape statistic “warrants maximal power for rejecting the [Yule] model against the PDA [model]”, this is hardly surprising. It can easily be explained by the fact that the $\widehat{s}$-shape statistic was even introduced with that very purpose in mind. In fact, except for the normalizing constant of $P_{Y,n}$, the $\widehat{s}$-shape statistic corresponds to the logarithm of $P_{Y,n}$ (c.f. [79]), i.e., of the probability of generating a tree under the Yule model. However, while Blum et al. in their study explicitly did not claim that the $\widehat{s}$-shape statistic is “generally superior”, in our study we could see that it is indeed superior concerning many models, not only the PDA model. Some exceptions include the Brownian models, in which the $\widehat{s}$-shape statistic is outperformed by other TSS, c.f. Figures 5 and 6. Khurana et al.  [42] also found that for their data, the $\widehat{s}$-shape statistic’s performance was mediocre, while the $B_{2}$ index, the variance of leaf depths, $I$-based indices, stairs2, and the normalized Colless index were most powerful in discriminating empirical trees from trees simulated under a constant-rate-birth-death model (which induces the same distribution on $\mathcal{BT}^{\ast}_{n}$ as the Yule model). In Section 6 of the supplementary material we discuss their findings further and show how to approach their research questions with our methods. We also remark that the findings by Khurana et al.  [42] concerning $B_{2}$ nicely complement a study by Bienvenu et al. [80], which used five different null models (one of which was the Yule model) and tested six different balance indices for their power (note that they did not include the $\widehat{s}$-shape statistic in their study, though). In this study, the authors found that the $B_{2}$-index is on average better than the other investigated indices. Note that in most of our studies, too, the $B_{2}$ index was amongst the most powerful indices.
@@
As shown by Bienvenu et al. [80], the choice of different null models can make a huge difference. However, as our results show, even with the Yule model as a null model, there are huge performance differences between the different TSS. These differences might imply that there are different patterns of imbalance produced by the different models, some of which can only be recognized by a subset of tree shape statistics. Therefore, the choice of the most powerful TSS for each individual case is important. We thus highly recommend to perform a power analysis such as the one presented here prior to downstream analyses. For this purpose, our software package poweRbal is specifically designed in such a way that it also caters to researchers unfamiliar with the programming language R. Trees generated by other software tools can easily be loaded into our package (e.g., in the well-known Newick format) and used as a basis for an alternative model. Newly developed indices and models can easily be added to the package, so that the user can test several variations to see which one is most powerful in the given setting. poweRbal also enables researchers to perform a simulation study beforehand to decide on which TSS to use and to subsequently analyze their data with the respective optimal tools.
@@
Concerning theoretical research, on the other hand, a goal for future developments could be to develop tools to guess the most likely model simply based on a given tree. First results in this regard were presented in [75], where the likelihood of pitchforks and cherries under Ford’s $\alpha$-model has been calculated, which enables users to find the most likely $\alpha$-value. However, as the possibilities of using pitchforks and cherries for estimating the best-fitting overall model are limited, more research in this regard is necessary.
@@
Finally, our manuscript has focused on basic models, i.e., on models which contain clear instructions on how to generate (phylogenetic) trees. However, in the literature also more elaborate and complex models can be found (e.g., [41, pp. 9–13] and [81]). For instance, evolution can be modeled at the genome level, so that subsequently a phylogenetic tree can be reconstructed (e.g., [81]). For the latter, different reconstruction methods can be used, and it has been shown that these methods also have an impact on the resulting tree’s balance (see, e.g., [82, 83], [84, p. 1141])). Moreover, we have primarily focused on models from phylogenetics, but tree models also frequently occur in other areas such as population genetics, probability theory, or computer science. While some of the models used in these areas are closely related to the Yule model (see, e.g., [64] and Remark 3.1 in the supplement of the present manuscript), it would be an interesting direction of future research to include models such as Aldous’ continuum random trees  [85, 86, 87] in the analyses.
@@ Acknowledgement
MF and SK were supported by the project ArtIGROW, which is a part of the WIR!-Alliance “ArtIFARM – Artificial Intelligence in Farming”, and gratefully acknowledge the Bundesministerium für Bildung und Forschung (Federal Ministry of Education and Research, FKZ: 03WIR4805) for financial support. Moreover, we wish to thank Volkmar Liebscher for helpful advice on the development of the R-package and Annalena Franke for implementing first versions of several tree models. Additionally, we thank three reviewers and the editor Noah Rosenberg for detailed comments on an earlier version of this paper.

### v4

@@ 4 Results
In this section, we summarize the most important results obtained in our simulation study. Figures 37 show the power (with 95%-confidence bands) of the various TSS to correctly identify trees generated under models distinct from the Yule model M-0 as not having been generated under the Yule model. The power is plotted either against the leaf number $n$ (Figure 3) or against a tree model parameter (Figures 47). Note that all figures follow the legend of Figure 3. Further result figures can be found in Section 5 of the supplementary material [40]. As evident from all our figures, the curves look quite different for different alternative models. In the following, we study some aspects of these figures and conclusions to be drawn more in-depth.
The most and least powerful TSS
@@
A central goal of our study was to identify TSS that exhibit a high power for discriminating between different (macro)evolutionary models. Our results show that the performance of different TSS clearly depends on the alternative models and also on the null model used (recall that we use the Yule model M-0 as the null model throughout the present manuscript, but in Section 6 of the supplementary material [40] we also briefly discuss other null models). Even for a fixed alternative model, there are sometimes striking differences for different parameter choices. For example, considering the log-Brownian models M-9 in Figure 6 (a) & (b), we can observe that the $B_{1}$ index I-1 and the equal weights Colless index I-12 are among the TSS with the lowest power for small values of $\sigma_{\lambda}$, whereas they are among the most powerful TSS for larger values of $\sigma_{\lambda}$. A related observation is that in some cases, there are few to no crossings between the power curves for different TSS (e.g., Figures 4 (a) and 5 (b)), implying that the ranking of the TSS by their power does not change across the different settings tested, whereas for others, there are significant changes (e.g., Figures 5 (a) and 6 (a) & (b)).
@@
Nevertheless, we can observe that there are some TSS that (almost) consistently outperform the others, such as the $\widehat{s}$-shape statistic I-17, whereas others, such as the modified maximum difference in widths I-6 or the rooted quartet index I-7, tend to consistently exhibit a low power. In all cases, there are notable exceptions, though. In the case of the $\widehat{s}$-shape statistic I-17 it is interesting to note that it is outperformed by other TSS (such as the $B_{2}$ index I-2 or the Sackin index I-16) in the symmetric simple Brownian model M-7 (Figure 5 (b)) as well as for parameter-model combinations close to the Yule model M-0 in the asymmetric log-Brownian model M-9 (Figure 6 (b))). In the case of the modified maximum difference in widths I-6, exceptions to its overall relatively low power are for instance Aldous’ $\beta$ splitting model M-1 (Figure 4 (a)) and the biased speciation model M-5 (Figure 5 (a)). Another exception for the rooted quartet index I-7 is the simple Brownian model M-7 (Figure 5 (b)). It is also noticeable that while the power of TSS should increase with increasing $n$, and this is indeed the case for almost all TSS (see Figure  3 and the differences between $n=30$ and $n=100$ in Figure 7 (a) & (b)), the power of the modified maximum difference in widths I-6 does not seem to follow this pattern for the PDA model M-15 (Figure 3).
@@
Finally, we remark that it is often the case that several TSS have equal power, indicated by their curves being clumped together, as is for instance the case for the top curves in the simple Brownian model M-7 (Figure 5 (b)). In other cases, the curves are clearly separated, as for instance in the biased speciation model M-5 (Figure 5 (a)). Interestingly, the total cophenetic index I-19 and the quadratic Colless index I-11 show extremely similar behavior (their lines are nearly always right on top of each other), and it would be interesting to investigate this further. In summary, we thus reiterate that which TSS have a high power and which have a low power depends on the model setting, highlighting the importance of performing a power analysis such as the one presented here prior to downstream analyses.
@@
For parameter settings that correspond to the Yule model M-0, the power of all TSS should be close to 0.05 and this is clearly visible in the corresponding figures (e.g., for $\beta=0$ or $\alpha=0$ in Figure 4 (a) & (b), or $\mu=0$ in Figure 7 (a) & (b)). Interestingly, we also observe a power close to 0.05 for almost all TSS for the biased speciation model M-5 for $\zeta$ between 0.20 and 0.25 (Figure 5 (a)). This could indicate that the biased speciation model M-5 is very close to or coincides with the Yule model M-0 for $\zeta$ in this range. However, it could also mean that none of the TSS employed in this study can detect deviation from the Yule model M-0 in this setting, even though the models might theoretically not be identical. This leads us to the following observation for fitting tree models to empirical data: When a collection of empirical trees is tested against a specific evolutionary model and a low power is observed for all TSS, this can be an indication of a good fit of the model to the data. However, low power is only a sufficient and not a necessary condition for a good model fit, as it could be the case that other TSS are needed to detect differences between the data and the model. Thus, caution is warranted. A good example illustrating this can be found in Figure 20 in the supplementary material [40], where we use Aldous’ $\beta$ splitting model M-1 with $\beta=-1$ as the null model and the alternative birth-death model as the alternative model. In this case, almost none of the TSS can tell the alternative birth-death trees with parameter $\mu=0.4$ apart from the Aldous $\beta=-1$ trees, except for the $B_{2}$ index I-2 and the Furnas rank I-3 (where the latter is even more surprising as it is primarily used as a tree enumeration method and not often employed in applications). Without including these two indices, one might have concluded that both models are equivalent.
@@
Figure 3: The power of all TSS to correctly identify trees generated under the PDA model M-15 as not having been generated under the Yule model M-0 as a function of $n$. Note that for $n=10$ and $n=15$ the power values are based on the exact TSS distribution under Yule. Notice that the legend of this figure also applies to all subsequent figures.
Figure 4: The power of all TSS to correctly identify trees generated under (a) the Aldous $\beta$ splitting model M-1 and (b) Ford’s $\alpha$ model M-14 (right) as not having been generated under the Yule model M-0 as a function of $\beta$ (resp. $\alpha$). Since for $\beta=0$, Aldous’ splitting model corresponds M-1 to the Yule model M-0, all TSS rejected $\approx 0.05$ of the trees in that case (as specified with the level of significance). Similarly, since for $\alpha=0$, Ford’s $\alpha$ model M-14 corresponds to the Yule model M-0, all TSS rejected $\approx 0.05$ of the trees in that case. Notice that this figure follows the same legend as Figure 3.
Figure 5: The power of all TSS to correctly identify trees generated under (a) the biased speciation model M-5 and (b) the asymmetric simple Brownian model M-7 as not having been generated under the Yule model M-0 as a function of $\zeta$ (resp. $\sigma$). Notice that this figure follows the same legend as Figure 3.
Figure 6: The power of all TSS to correctly identify trees generated under (a) the symmetric and (b) the asymmetric punctuated(-intermittent) log-Brownian model M-9 as not having been generated under the Yule model M-0 as a function of $\sigma_{\lambda}$. The starting trait value is $x_{0}=10$ with constant $\sigma_{x}=1$. Notice that this figure follows the same legend as Figure 3.
Figure 7: The power of all TSS to correctly identify trees generated under the alternative birth-death model M-11 as not having been generated under the Yule model M-0 as a function of $\mu$ for two different tree sizes ((a) $n=30$ and (b) $n=100$). Notice that this figure follows the same legend as Figure 3.
@@ 5 Conclusion and outlook
In most of our analyses, the $\widehat{s}$-shape statistic I-17 was the most powerful statistic discriminating alternative models from the Yule model M-0. While this nicely complements a study by Blum at al. [27], who found that the $\widehat{s}$-shape statistic I-17 “warrants maximal power for rejecting the [Yule] model against the PDA [model]”, this is hardly surprising. It can easily be explained by the fact that the $\widehat{s}$-shape statistic I-17 was even introduced with that very purpose in mind. In fact, except for the normalizing constant of $P_{Y,n}$, the $\widehat{s}$-shape statistic I-17 corresponds to the logarithm of $P_{Y,n}$ (c.f. [80]), i.e., of the probability of generating a tree under the Yule model M-0. However, while Blum et al. in their study explicitly did not claim that the $\widehat{s}$-shape statistic I-17 is “generally superior”, in our study we could see that it is indeed superior concerning many models, not only the PDA model M-15. Some exceptions include the Brownian models M-7 and M-9, in which the $\widehat{s}$-shape statistic I-17 is outperformed by other TSS, cf. Figures 5 (b) and 6 (b). Khurana et al.  [42] also found that for their data, the $\widehat{s}$-shape statistic I-17 had relatively low power, whereas the $B_{2}$ index I-2, the variance of leaf depths I-21, the mean and total $I$ indices I-14 and I-20 (without correction method $I^{\prime}$), stairs2 I-8, and the normalized (corrected) Colless index (I-9) were most powerful in discriminating empirical trees from trees simulated under a constant-rate-birth-death model (which induces the same distribution on $\mathcal{BT}^{\ast}_{n}$ as the Yule model M-0). In Section 6 of the supplementary material we discuss their findings further and show how to approach their research questions with our methods. We also remark that the findings by Khurana et al.  [42] concerning the $B_{2}$ index I-2 nicely complement a study by Bienvenu et al. [81], which used five different null models (one of which was the Yule model M-0) and tested six different balance indices for their power (note that they did not include the $\widehat{s}$-shape statistic I-17 in their study, though). In this study, the authors found that the $B_{2}$-index I-2 has on average higher power than the other investigated indices. Note that in most of our studies, too, the $B_{2}$ index I-2 was amongst the most powerful indices.
@@
As shown by Bienvenu et al. [81], the choice of different null models can make a huge difference. However, as our results show, even with the Yule model M-0 as a null model, there are huge performance differences between the different TSS. These differences might imply that there are different patterns of imbalance produced by the different models, some of which can only be recognized by a subset of tree shape statistics. Therefore, the choice of the most powerful TSS for each individual case is important. We thus highly recommend to perform a power analysis such as the one presented here prior to downstream analyses. For this purpose, our software package poweRbal is specifically designed in such a way that it also caters to researchers unfamiliar with the programming language R. Trees generated by other software tools can easily be loaded into our package (e.g., in the well-known Newick format) and used as a basis for an alternative model. Newly developed indices and models can easily be added to the package, so that the user can test several variations to see which one is most powerful in the given setting. poweRbal also enables researchers to perform a simulation study beforehand to decide on which TSS to use and to subsequently analyze their data with the respective optimal tools.
@@
Concerning theoretical research, on the other hand, a goal for future developments could be to develop tools to guess the most likely model simply based on a given tree. First results in this regard were presented in [77], where the likelihood of pitchforks and cherries under Ford’s $\alpha$-model M-14 has been calculated, which enables users to find the most likely $\alpha$-value. However, as the possibilities of using pitchforks and cherries for estimating the best-fitting overall model are limited, more research in this regard is necessary.
Another interesting direction for future research would be to investigate whether there is a connection between the power and resolution of different TSS. Specifically, it is interesting to note that Fischer et al. [82, Chapter 26] recently observed that the $\widehat{s}$-shape statistic I-17 had a high resolution for two empirical data sets (meaning that it showed few ties), whereas the modified maximum difference in widths I-6, an index we observed to have a relatively low power in a variety of settings, showed a low resolution (meaning that it showed many ties). This could indicate a connection between resolution and power for some but not all TSS. For the rooted quartet index I-7, for instance, Fischer et al. [82, Chapter 26] observed a high resolution while we observe a relatively low power of this index in the present analysis. Nevertheless, exploring why certain TSS behave in distinct ways and whether there is a connection between power and resolution for some TSS is an interesting direction for future research.
@@
Finally, our manuscript has focused on basic models, i.e., on models which contain clear instructions on how to generate (phylogenetic) trees. However, in the literature also more elaborate and complex models can be found (e.g., [41, pp. 9–13] and [83]). For instance, evolution can be modeled at the genome level, so that subsequently a phylogenetic tree can be reconstructed (e.g., [83]). For the latter, different reconstruction methods can be used, and it has been shown that these methods also have an impact on the resulting tree’s balance (see, e.g., [84, 85], [86, p. 1141]). Moreover, we have primarily focused on models from phylogenetics, but tree models also frequently occur in other areas such as population genetics, probability theory, or computer science. While some of the models used in these areas are closely related to the Yule model M-0 (see, e.g., [66] and Remark 3.1 in the supplement [40] of the present manuscript), it would be an interesting direction of future research to include models such as Aldous’ continuum random trees  [87, 88, 89] in the analyses.
@@ Acknowledgement
MF and SK were supported by the project ArtIGROW, which is a part of the WIR!-Alliance “ArtIFARM – Artificial Intelligence in Farming”, and gratefully acknowledge the Bundesministerium für Bildung und Forschung (Federal Ministry of Education and Research, FKZ: 03WIR4805) for financial support. Moreover, we wish to thank Volkmar Liebscher for helpful advice on the development of the R package and Annalena Franke for implementing first versions of several tree models. Additionally, we thank three reviewers and the editor Noah Rosenberg for detailed comments on an earlier version of this paper.